OpenStudy (anonymous):

There are six numbers on a normal die. What is the probability you will roll an odd number? Reduce the fraction to its simplest form. 1/6 1/4 1/3 1/2

jimthompson5910 (jim_thompson5910):

how many odd numbers are there on a normal die?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

are you sure? here are the list of numbers on the die 1, 2, 3, 4, 5, 6

OpenStudy (anonymous):

oh my bad 3

OpenStudy (anonymous):

so its 1/3

jimthompson5910 (jim_thompson5910):

there are 3 odd numbers out of 6 total

jimthompson5910 (jim_thompson5910):

so, 3/6 = ???

OpenStudy (anonymous):

0.5

jimthompson5910 (jim_thompson5910):

which is what as a fraction?

OpenStudy (anonymous):

1/3

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

to reduce 3/6, you just divide both parts by 3

jimthompson5910 (jim_thompson5910):

since 3 is the GCF

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so i divide them both and i got 1/2

jimthompson5910 (jim_thompson5910):

good, 3/6 = 1/2